OpenStudy (12wishs):

Can I get some help please? What is the product of three square root of eight times four square root of three? Simplify your answer. Twenty four square root six Twelve square root twelve Twelve square root eleven Seventy two square root four

OpenStudy (phi):

do you mean \[ 3 \sqrt{8} \cdot 4 \sqrt{3} \]

OpenStudy (alexandervonhumboldt2):

\[\huge\bf 3\sqrt{8} * 4\sqrt{3}?\] does it looks like tihs?

OpenStudy (12wishs):

yes

OpenStudy (alexandervonhumboldt2):

we combine 3 with 4 and sqrt8 with sqrt3: \[\large\rm so~we~have~~(3*4)*\sqrt{8*3}\]

OpenStudy (12wishs):

ok

OpenStudy (alexandervonhumboldt2):

simplify this

OpenStudy (12wishs):

29.39

OpenStudy (phi):

when you do "square root" problems, it is usually helpful to factor the numbers inside the square root for example, write 8 as 2*2*2 and 8*3 as 2*2*2*3 and then look for "pairs" of numbers to "pull out"

OpenStudy (alexandervonhumboldt2):

no, write this in sqrt form. do not convert it into decimal

OpenStudy (12wishs):

how?

OpenStudy (alexandervonhumboldt2):

well what is 8*3?

OpenStudy (12wishs):

24

jhonyy9 (jhonyy9):

@12wishs so there is 3sqrt8 than you know that sqrt8 = sqrt(4*2) but you know that srt(4) = 2 so than how can you rewriting 3sqrt8 = ?

OpenStudy (12wishs):

2^3?

jhonyy9 (jhonyy9):

sqrt8 = sqrt(2^2 *2) = ?

OpenStudy (12wishs):

8

jhonyy9 (jhonyy9):

for example sqrt12 = sqrt(4*3) = sqrt(2^2 *3) = 2sqrt3 us this in case of your exercise please

OpenStudy (mathmale):

There are several ways in which you could simplify "\[3 \sqrt{8} \cdot 4 \sqrt{3}\]

OpenStudy (mathmale):

...and at least two have been mentioned by others. Method 1: You can combine the two radicals Sqrt(8) and Sqrt(3) into a single radical:\[\sqrt{8}\sqrt{3}\sqrt{24}\]

OpenStudy (mathmale):

That's = to Sqrt(24), sorry.

jhonyy9 (jhonyy9):

@mathmale there you have missed a sign of =

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (mathmale):

Next, in this same example, we can factor 24 into one factor that IS a perfect square and one factor which is not. Could you do that, please?

OpenStudy (12wishs):

4 is a perfect square and can be considered the square root of 24 and 3 can be considered the square root of 24 and it is not a perfect square

OpenStudy (mathmale):

Method 2: starting from \[3 \sqrt{8} \cdot 4 \sqrt{3}\]

OpenStudy (mathmale):

... factor that 8 in the same way, again identifying the one perfect square factor and the one non-perfect-square factor. Could you do that, please?

OpenStudy (mathmale):

@12wishs ??

OpenStudy (12wishs):

Perfect square=4 and imperfect square=3

OpenStudy (mathmale):

Great. Now, what's the square root of 4? of 3?

OpenStudy (12wishs):

the square root of 4 is 2 and the square root of 3 is 3

OpenStudy (mathmale):

Except that you must write \[\sqrt{3}\]

OpenStudy (mathmale):

Also note that Sqrt(24)=Sqrt(4)*Sqrt(6).

OpenStudy (mathmale):

\[\sqrt{24}=\sqrt{4}\sqrt{6}\]

OpenStudy (mathmale):

or 2\[\sqrt{6}\]

OpenStudy (mathmale):

Could you use this result to simplify\[3 \sqrt{8} \cdot 4 \sqrt{3}\]

OpenStudy (mathmale):

Note that \[3 \sqrt{8} \cdot 4 \sqrt{3}\]

OpenStudy (mathmale):

... can be rearranged as\[24\sqrt{24}=24*2*\sqrt{6}\]

OpenStudy (mathmale):

How'd you simplify that?

OpenStudy (12wishs):

simplify 6 to its square root of 2 then multiply 24 by 2 which would be 48 then again by 2 which would be 96

OpenStudy (mathmale):

Hint: 24*2=?

jhonyy9 (jhonyy9):

@mathmale there is something wrong please check it

jhonyy9 (jhonyy9):

12sqrt24 not 24sqrt24 i think

OpenStudy (mathmale):

I agree that I've stretched this discussion 'way out (without meaning to), and that the problem could be done much faster. However, we end up with 48 times what? 12wishs?

OpenStudy (12wishs):

so wait would the answer to the question be 24 sqrt6?

OpenStudy (mathmale):

Yes, indeed. So sorry this has taken so long.

OpenStudy (12wishs):

its ok I'm just glad I got an answer

OpenStudy (12wishs):

thanks for your help

jhonyy9 (jhonyy9):

GREAT @mathmale !!! hope so much that @12wishs have understood it all right now sure

OpenStudy (mathmale):

You're welcome. My face is red with embarrassment, however.

OpenStudy (mathmale):

;(